OpenStudy (anonymous):

Convert the given exponential function to the form indicated. f(t) = 3.1(1.007)^t; f(t) = Q0e^kt

OpenStudy (anonymous):

\[\huge f(t)= 3.1\left(1+\frac{1}{\frac{1000}{7}} \right)^{\frac{1000}{7}t} = 3.1e^{\frac{1000}{7}t} \]

OpenStudy (anonymous):

there